OpenStudy (anonymous):

A 30-day supply of vitamins costs $2.91, what is the price in cents per day? The vitamins cost cents per day

OpenStudy (acxbox22):

divide cost by number of days since we are finding cents per day

OpenStudy (texaschic101):

30/291 = 1/x --- (30 days supply to 291 cents = 1 day supply to x cents) this is a proportion, so we cross multiply 30x = 291 x = 291/30 x = ??....you can do the math